Romeo is 3 yrs younger than his sister khezel,6 yrs hence,khezel will be thrice as old as romeo was five years ago.how old is romeo?
Answer by lwsshak3(11628)

Romeo is 3 yrs younger than his sister khezel,6 yrs hence,khezel will be thrice as old as romeo was five years ago.how old is romeo?
***
let x=Romeo's present age
x+3=Sister khezel's present age
..
6 yrs hence:
x+3+6=x+9=khezel's age
..
5 yrs ago:
x-5=Romeo's age
..
x+9=3(x-5)
x+9=3x-15
2x=24
x=12
how old is romeo? 12